7) Sixty-four percent of an axolotl breeding colony are albinos. In axolotls, albinism has a classic dominant-recessive inheritance pattern with wildtype being dominant to albino. What are the frequencies of the dominant and recessive alleles in this axolotl colony? Also, what percent of the population is heterozygous? Show your work!
Answer #1

Here, 64% are albinos. So, (100-64) = 36% are wild type.

Suppose, frequency of albino allele is p and wild type allele is q.

Here, p^2 = 64/100 = 0.64

So, p = 0.8

According to Hardy Weinberg law,

p + q = 1

Now, q = 1-0.8 = 0.2

So, frequencies of the dominant and recessive alleles in this axolotl colony are 0.2 and 0.8.

Now, according to Hardy Weinberg law,

p^2 + 2pq + q^2 = 1

Here, we are interested to know the heterozygous, that is 2pq.

It is 2 X 0.8 X 0.2 = 0.32

So, the percentage of the heterozygous population is 0.32 X 100 = 32%.